Teacher's own problem, not loacted in a book. Three consecutive even integers are such that the square of the third is 76 more than the square of the second. Find the three integers.
Answer by Prithwis(166)

Let the consecutive even integers be (x-2),x,(x+2)
(x+2)^2 = 76 + x^2
=> x^2+4x+4 = x^2+76
=> x = 18
The consecutive integers are 16, 18, 20